OpenStudy (anonymous):

(I know its another log question, im obviously not very good at these ha) evaluate the expression: log8(subscript)64^7.

OpenStudy (anonymous):

\[\log_ab^c = c\log_ab \] and \[log_xx^2 = 2 \]

OpenStudy (anonymous):

ans is 1/14

OpenStudy (anonymous):

how to write these mathematical expressions

OpenStudy (anonymous):

By clicking equation (fun fact: it's not 1/14)

OpenStudy (anonymous):

so would it be 7log8? since 64/8 is 8?

OpenStudy (anonymous):

logx(subscript) \[x ^{m}\] then answer is 1/m sorry guys i'm not good at typing mathematical equations

OpenStudy (anonymous):

and i'm sure it ans is 1/14

OpenStudy (anonymous):

\[\log_864^7 \not= \frac{1}{14} \]

OpenStudy (anonymous):

was i correct? possibly? maybe? ha

OpenStudy (anonymous):

if m is the power to subscript then it comes as 1/m before log

OpenStudy (anonymous):

hey inewton \[(64)^{7}\] is subscript

OpenStudy (anonymous):

No it isn't, because that makes no sense.

OpenStudy (anonymous):

the 8 is the subscript

OpenStudy (anonymous):

the 64 is the base, and the 7 is the superscript

OpenStudy (anonymous):

ans is 1/14. question is \[\log_{64^{7}} \]8

OpenStudy (anonymous):

i mean log8 (subcript)\[64^{7}\]

OpenStudy (anonymous):

the question is \[\log_864^7 = 7\log_864 \]

OpenStudy (anonymous):

ok, thank you very much

OpenStudy (anonymous):

But log_8 64 simplifies further

OpenStudy (anonymous):

it simplifies to just 8, right?

OpenStudy (anonymous):

No \[\log_ab = c \iff a^c = b\] So log_8 64 = 2 because 8^2 = 64

OpenStudy (anonymous):

ohhh, ok. thank you! sorry it took so long!

OpenStudy (anonymous):

You're welcome